# MeshAnything: Artist-Created Mesh Generation with Autoregressive Transformers

### v1

@@ 4.1 Shape Encoding for Conditional Generation
    It should be easily extracted from various 3D representations. This ensures that the trained models can be integrated with a wide range of 3D asset production pipelines [44, 34, 30, 47, 57].
@@
Therefore, we need to ensure that $\mathcal{S}$ extracted from the ground truth $\mathcal{M}$ during training has a similar domain to the $\mathcal{S}$ extracted during inference. To bring their domains closer, we intentionally construct coarse meshes from AMs. We first extract the signed distance function from $\mathcal{M}$ with [63], then convert it into a relatively coarse mesh using Marching Cubes [42] to destroy the ground truth topology. Finally, we sample point cloud and its normals from the coarse mesh. This approach also helps to avoid overfitting, as AMs typically have fewer faces, and each face can often sample multiple points. The network can easily recognize the ground truth topology by determining whether the points lie on the same plane.
Since almost all 3D representations can be converted into a coarse mesh using Marching Cubes [42] or sampled into point clouds, this ensures that the domain of $\mathcal{S}$ is consistent during both training and inference. We pair the point clouds extracted as $\mathcal{S}$ with $\mathcal{M}$ to create a data item $\{(\mathcal{M}_{i},\mathcal{S}_{i})\}_{i}$ for training.
@@ 4.2 VQ-VAE with Noise-Resistant Decoder
Following MeshGPT [55], we first train a VQ-VAE [61] to learn a vocabulary of geometric embeddings for better transformer [62] learning. Different to MeshGPT, which uses graph convolutional networks [67] and ResNet [29] as the encoder and decoder respectively, we employ transformers with identical structures for both the encoder and decoder. When training VQ-VAE, meshes are discretized and input as a sequence of triangle faces:
@@
The VQ-VAE is trained end-to-end with cross-entropy loss on the predicted vertex coordinate logits and the commitment loss of vector quantization [61]. After the training of VQ-VAE, the encoder-decoder of VQ-VAE is treated as a tokenizer and detokenizer for autoregressive transformer training.
@@ 4.3 Shape-Conditioned Autoregressive Transformer
To add shape condition to the transformer, inspired by the success of multimodal large language models [68, 37, 70, 25], we first encode the point cloud into a fixed-length token sequence with a point cloud encoder $\mathcal{P}$ and then concatenate it to the front of the embedding sequence from $\mathcal{T}$ VQ-VAE as the final input embedding sequence for the transformer:
@@
We borrow a pretrained point encoder from [75] and add a linear projection layer to project its output feature to the same latent space as $\mathcal{T}$. During training, the original point encoder from [75] is frozen; we only update the newly added projection layer and the autoregressive transformer with cross-entropy loss.
@@
Table 1: Reconstruction Performance under Different Noise Levels with and without Noise-Resistant (NR) Decoder. Please refer to 5.4 for metrics explanation.
@@
^$\hat{\mathcal{M}}=D(\hat{\mathcal{T}})$ (10)
where $\hat{\mathcal{M}}$ is the final generated AM.
5 Experiments
5.1 Data Preparation
Data Selection. Existing AM generation works are limited to a few categories. However, our method targets to operate on general shapes. MeshAnything is trained on a combined dataset of Objaverse [19] and ShapeNet [7], selected for their complementary characteristics. We chose Objaverse because it contains a large number of AMs without category limitations. On the other hand, ShapeNet offers higher data quality within limited categories.
Data Processing and Augmentation. Following the strategies of PolyGen [45] and MeshGPT [55], we order faces by their lowest vertex index, then by the next lowest, and so on. Vertices are sorted in ascending order based on their z-y-x coordinates, where z represents the vertical axis. Within each face, we permute the indices to ensure the lowest index comes first. During training, we apply on-the-fly scaling, shifting, and rotation augmentations, normalizing each mesh to a unit bounding box from $-0.5$ to $0.5$.
5.2 Implementation Details
The encoder and decoder of VQ-VAE both use the encoder of BERT [20], while we choose OPT-350M [74] as our autoregressive transformer architecture. The residual vector quantization [73] depth is set to 3, with a codebook size of 8,192.
Our point encoder is based on the pretrained point encoder from [75], which has been trained on Objaverse and thus can handle general shapes. This point encoder outputs a fixed-length token sequence of 257 tokens, with 256 tokens primarily containing shape information and an additional head token containing semantic information about the shape. We sample 4096 points for each point cloud.
Table 2: Ablation on Noise-Resistant (NR) Decoder for the Quality of Mesh Generation. Please refer to 5.4 for metrics explanation.
@@ 5.2 Implementation Details
Table 3: Quantitative evaluation with Marching Cubes [42] and Remesh [4]. Please refer to 5.4 for metrics explanation.
@@
(b) Remesh (0.005) 2.174 7.813 0.912 127.8 167.9 748.1 534.6
(c) Remesh (0.010) 2.083 7.578 0.929 39.01 41.78 225.4 132.3
(d) Remesh (0.030) 2.915 8.329 0.863 5.848 4.410 34.38 14.05
(e) Remesh (0.050) 4.179 8.138 0.814 2.299 1.538 13.64 4.920
(f) Remesh (0.100) 7.312 10.771 0.748 0.625 0.359 3.735 1.149
(g) MeshAnything 2.256 6.245 0.902 0.172 0.318 0.888 0.871
The VQ-VAE is trained on 8 A100 GPUs for 12 hours and the transformer is trained on 8 A100 GPUs for 4 days.
5.3 Qualitative Experiments
We present more qualitative results of our model combined with other 3D asset production pipelines. As shown in the Fig. 1, Fig. 4 and Fig. 5, MeshAnything effectively generates AMs from various 3D representations. When integrated with different 3D assets production pipelines, our method effectively achieves mesh generation with diverse conditions.
Next, Fig. 5 demonstrates that our model does not simply overfit but understands how to generate meshes with efficient topology that conform to the given shape. To prove this, we use manually-created meshes as ground truth and use their shapes as conditions to test whether our model can generate meshes with comparable topology. To effectively use the ground truth as conditions, we first convert them into dense meshes using Marching Cubes [42] to disrupt their face structure. Then, we sample point clouds with normals from the dense meshes to serve as shape conditions. The experimental results in Fig. 5 show that MeshAnything is capable of generating meshes comparable to or even surpassing those modeled by human artists, exhibiting diverse and strong 3D modeling capabilities.
5.4 Quantitative Experiments
@@
Metrics. We follow the evaluation metric setting of [12]. We quantitatively evaluate mesh quality by uniformly sampling 100K points from the faces of both the ground truth meshes and the predicted meshes, and then computing a set of metrics to assess various aspects of the reconstruction. We report the following metrics: Chamfer Distance (CD) to evaluate the overall quality of a reconstructed mesh; Edge Chamfer Distance (ECD) to assess the preservation of sharp edges by sampling points near sharp edges and corners; and Normal Consistency (NC) to evaluate the quality of the surface normals. Additionally, we report the number of mesh vertices (#V) and the number of mesh faces (#F). We also provide the ratio of the estimated number of vertices to the ground truth number of vertices (#V_Ratio) and the same ratio for faces (#F_Ratio).
@@
Ablations on Noise-Resistant Conditional Decoder. We first perform ablation experiments to verify the effectiveness of the Noise-Resistant Decoder. We begin with a VQ-VAE trained without any noise or conditioning. We then perform ablation between two settings: one where the decoder remains unchanged and unaware of the shape condition, and another where the shape condition is injected into the transformer, as described in Section 4.2. Next, we randomly sample a noise from gumbel distribution and add it to codebook sampling logits during the vector quantization process to simulate the potential low-quality token sequences generated by the transformer. We control the noise level by scaling the added noise.
@@
Comparison with Marching Cubes and Remesh. Our method is related to various mesh extraction methods [42, 13, 12, 53] since we all convert other 3D representations into meshes. However, it is important to note that previous approaches are reconstruction-like methods that produce dense meshes, while our approach is generative, creating Artist-Created Meshes (AMs) that are significantly more complex to produce than dense meshes. Current metrics can only indicate the quality of shape alignment, which do not effectively reflect the topological advantages of our method. Among the numerous mesh extraction methods, we chose the most representative one, Marching Cubes [42], for comparison.
We compare the results of mesh extraction using Marching Cubes [42] and its combination with Blender remesh [4]. Since our evaluation dataset includes non-watertight meshes, we first extract the signed distance fields (SDF) of all ground truth meshes using [63], which can handle non-watertight meshes. We then apply Marching Cubes with a resolution of 128 on these SDFs. Next, we apply Blender remesh [4] with different voxel sizes to the Marching Cubes results, as both the remesh method and our approach are capable of simplifying topology. Additionally, the Marching Cubes result is used as the shape condition input to MeshAnything to obtain our results.
As shown in Tab. 3, we found that these methods require hundreds of times more faces to achieve results comparable to our method. Furthermore, our approach achieves the best performance in Edge Chamfer Distance (ECD), even with significantly fewer faces than other methods. Comparing (a) and (g), our method lags in Chamfer Distance (CD) and Normal Consistency (NC), mainly due to our method’s inherent failure cases as a generative model, which makes it less robust than Marching Cubes. When comparing with remesh methods, we observe that they incur a high cost to achieve a face count similar to ours. Comparing (f) and (g), we find that even when remesh methods achieve a comparable face count, the number of vertices is still several times higher than ours, indicating that the topology efficiency of remesh methods is far inferior to ours, as they completely ignore the shape characteristics of the 3D assets. Additionally, we surprisingly find that our method can produce results with fewer faces than the ground truth, demonstrating that MeshAnything is not overfitting to the data but instead learns an efficient topology representation, occasionally surpassing the ground truth meshes.
6 Limitations
Our method cannot generate meshes that exceed the maximum face count limit, so it cannot convert large scenes and particularly complex objects into meshes. Additionally, due to its generative nature, our method is not as stable as methods like Marching Cubes [42].
7 Social Impact

### v2

@@ 4.1 Shape Encoding for Conditional Generation
    It should be easily extracted from various 3D representations. This ensures that the trained models can be integrated with a wide range of 3D asset production pipelines Mildenhall et al. (2020); Kerbl et al. (2023b); Hong et al. (2023); Poole et al. (2023); Tang et al. (2024).
@@
Therefore, we need to ensure that $\mathcal{S}$ extracted from the ground truth $\mathcal{M}$ during training has a similar domain to the $\mathcal{S}$ extracted during inference. To bring their domains closer, we intentionally construct coarse meshes from AMs. We first extract the signed distance function from $\mathcal{M}$ with Wang et al. (2022), then convert it into a relatively coarse mesh using Marching Cubes Lorensen & Cline (1987) to destroy the ground truth topology. Finally, we sample point cloud and its normals from the coarse mesh. This approach also helps to avoid overfitting, as AMs typically have fewer faces, and each face can often sample multiple points. The network can easily recognize the ground truth topology by determining whether the points lie on the same plane.
Since almost all 3D representations can be converted into a coarse mesh using Marching Cubes Lorensen & Cline (1987) or sampled into point clouds, this ensures that the domain of $\mathcal{S}$ is consistent during both training and inference. We pair the point clouds extracted as $\mathcal{S}$ with $\mathcal{M}$ to create a data item $\{(\mathcal{M}_{i},\mathcal{S}_{i})\}_{i}$ for training.
@@ 4.2 VQ-VAE with Noise-Resistant Decoder
Following MeshGPT Siddiqui et al. (2023), we first train a VQ-VAE Van Den Oord et al. (2017) to learn a vocabulary of geometric embeddings for better transformer Vaswani et al. (2017) learning. Different to MeshGPT, which uses graph convolutional networks Wu et al. (2019) and ResNet He et al. (2016) as the encoder and decoder respectively, we employ transformers with identical structures for both the encoder and decoder. When training VQ-VAE, meshes are discretized and input as a sequence of triangle faces:
@@
The VQ-VAE is trained end-to-end with cross-entropy loss on the predicted vertex coordinate logits and the commitment loss of vector quantization Van Den Oord et al. (2017). After the training of VQ-VAE, the encoder-decoder of VQ-VAE is treated as a tokenizer and detokenizer for autoregressive transformer training.
@@ 4.3 Shape-Conditioned Autoregressive Transformer
To add shape condition to the transformer, inspired by the success of multimodal large language models Wu et al. (2023); Liu et al. (2024a); Xu et al. (2023); Guo et al. (2023), we first encode the point cloud into a fixed-length token sequence with a point cloud encoder $\mathcal{P}$ and then concatenate it to the front of the embedding sequence from $\mathcal{T}$ VQ-VAE as the final input embedding sequence for the transformer:
@@
We borrow a pretrained point encoder from Zhao et al. (2024) and add a linear projection layer to project its output feature to the same latent space as $\mathcal{T}$. During training, the original point encoder from Zhao et al. (2024) is frozen; we only update the newly added projection layer and the autoregressive transformer with cross-entropy loss.
Table 1: Comparison of Mesh Generation Methods. As shown in the left table, compared to the baseline Artist-Created Mesh Generation method, the meshes generated by MeshAnything are better aligned with human preferences. In the right table, we compare MeshAnything with mesh extraction baselines, and it received the most votes. For detailed settings, please refer to Section 5.4.
[…]
PolyGen 12.7% 11.1%
MeshGPT 24.1% 28.2%
MeshAnything 63.2% 60.7%
[…]
MarchingCubes 38.1% 10.2%
Shape As Points 17.3% 6.2%
MeshAnything 44.6% 83.6%
^$\hat{\mathcal{M}}=D(\hat{\mathcal{T}})$ (10)
where $\hat{\mathcal{M}}$ is the final generated AM.
5 Experiments
5.1 Data Preparation
Data Selection. Existing AM generation works are limited to a few categories. However, our method targets to operate on general shapes. MeshAnything is trained on a combined dataset of Objaverse Deitke et al. (2023b) and ShapeNet Chang et al. (2015), selected for their complementary characteristics. We chose Objaverse because it contains a large number of AMs without category limitations. On the other hand, ShapeNet offers higher data quality within limited categories.
Data Processing and Augmentation. Following the strategies of PolyGen Nash et al. (2020) and MeshGPT Siddiqui et al. (2023), we order faces by their lowest vertex index, then by the next lowest, and so on. Vertices are sorted in ascending order based on their z-y-x coordinates, where z represents the vertical axis. Within each face, we permute the indices to ensure the lowest index comes first. During training, we apply on-the-fly scaling, shifting, and rotation augmentations, normalizing each mesh to a unit bounding box from $-0.5$ to $0.5$.
5.2 Implementation Details
The encoder and decoder of VQ-VAE both use the encoder of BERT Devlin et al. (2018), while we choose OPT-350M Zhang et al. (2022) as our autoregressive transformer architecture. The residual vector quantization Zeghidour et al. (2021) depth is set to 3, with a codebook size of 8,192.
Our point encoder is based on the pretrained point encoder from Zhao et al. (2024), which has been trained on Objaverse and thus can handle general shapes. This point encoder outputs a fixed-length token sequence of 257 tokens, with 256 tokens primarily containing shape information and an additional head token containing semantic information about the shape. We sample 4096 points for each point cloud.
The training batch size for both the VQ-VAE and the transformer is set to 8 per GPU. The VQ-VAE is trained on 8 A100 GPUs for 12 hours, after which we separately finetune the decoder part of the VQ-VAE into a noise-resistant decoder, as detailed in Section 4.2. Following this, the transformer is trained on 8 A100 GPUs for 4 days.
Table 2: Quantitative Comparisons with Prior Arts on Objaverse. MeshAnything significantly outperforms prior methods across all metrics. MMD, KID are scaled by $10^{3}$.
[…]
PolyGen 23.2 6.22 88.2 48.8 27.7
MeshGPT 41.7 3.83 67.3 25.1 6.11
MeshAnything 53.1 2.72 55.7 14.5 1.89
5.3 Qualitative Experiments
As shown in Fig. 1, MeshAnything effectively generates AMs from various 3D representations. In our experiments, we use Rodin Team (2024) as the text-to-3D and image-to-3D method, and employ Mildenhall et al. (2020) and Kerbl et al. (2023a) as the 3D reconstruction pipeline to obtain the corresponding NeRF and Gaussian Splatting models. For additional qualitative results, please refer to  A.2 combined with other 3D asset production pipelines.
5.4 Quantitative Experiments
From the generative model perspective, MeshAnything is a shape-conditioned mesh generation model. From the mesh extraction perspective, it extracts artist-created meshes from point clouds. Consequently, we compare MeshAnything with both types of methods. Additional experiments can be found in Appendix Section A.2.
User Study. As shown in Tab. 1, we conducted two user studies, comparing with mesh generation baselines Nash et al. (2020); Siddiqui et al. (2023) and mesh extraction baselines Lorensen & Cline (1987); Peng et al. (2021), respectively. The mesh generation baselines are trained on ShapeNet, and to ensure a fair comparison, we retrained them on Objaverse using the same transformer model as MeshAnything. Since the mesh generation baselines are all unconditional mesh generation methods, whereas MeshAnything is a shape-conditioned mesh generation method, we sampled shapes randomly from the evaluation set of Objaverse as inputs for MeshAnything, while for the baseline methods, we performed random sampling directly.
In the mesh extraction baseline, since our method can also be viewed as a point cloud to mesh approach, we included Peng et al. (2021), a point cloud to mesh method, as a baseline. Additionally, we optimized the results from the mesh extraction baseline using the Blender remesh method Blender Development Team (2024) to simplify the topology.
We collected $30$ results from each method and asked users to vote for the best one in terms of shape quality and topology quality. A total of $41$ users participated, providing 1,230 valid comparisons. Both user studies demonstrated the superiority of our method. The only difference between the retrained MeshGPT and MeshAnything is whether they are shape-conditioned, further proving the advantages of the shape-conditioned mesh generation setting.
Metrics. We follow the metric setting of Chen et al. (2022); Siddiqui et al. (2023). We detail this setting in Appendix Section. A.1.
Comparison with Mesh Generation Pipelines. We use the same retrained models from the user study for comparison. As shown in Tab. 2, MeshAnything significantly outperforms prior methods Nash et al. (2020); Siddiqui et al. (2023), indicating that it’s superior in both the shape and topology quality. Since the only difference between the retrained MeshGPT and MeshAnything is the inclusion of shape conditioning, the superior performance of MeshAnything further demonstrates that Shape-Conditioned Mesh Generation is a more suitable setting for mesh generation.
6 Conclusion
Appendix A Appendix
@@
Table 3: Reconstruction Performance under Different Noise Levels with and without Noise-Resistant (NR) Decoder. Please refer to A.1 for metrics explanation.
@@
Table 4: Ablation on Noise-Resistant (NR) Decoder for the Quality of Mesh Generation.
@@
Table 5: Quantitative evaluation with mesh extraction baselines. MC, FC, SAP refer to Marching Cubes Lorensen & Cline (1987), FlexiCubes Shen et al. (2023), and Shape As Points Peng et al. (2021), respectively. Please refer to A.1 for metrics explanation.
@@
(b) MC+Remesh (0.005) 2.174 7.813 0.912 127.8 167.9 748.1 534.6
(c) MC+Remesh (0.010) 2.083 7.578 0.929 39.01 41.78 225.4 132.3
(d) MC+Remesh (0.030) 2.915 8.329 0.863 5.848 4.410 34.38 14.05
(e) MC+Remesh (0.050) 4.179 8.138 0.814 2.299 1.538 13.64 4.920
(f) MC+Remesh (0.100) 7.312 10.771 0.748 0.625 0.359 3.735 1.149
(g) FC 1.190 6.121 0.967 59.12 121.1 378.2 391.1
(h) FC+Remesh (0.010) 1.861 6.940 0.933 37.98 40.19 205.5 124.2
(i) SAP 1.771 7.112 0.939 79.12 152.3 481.2 489.3
(j) SAP+Remesh (0.010) 2.367 7.862 0.925 39.17 42.87 239.1 136.6
(k) MeshAnything 2.256 6.245 0.902 0.172 0.318 0.888 0.871
A.1 Metrics
We follow the evaluation metric setting of Siddiqui et al. (2023) in mesh generation experiments and the setting of Chen et al. (2022) in mesh extraction experiments.
@@ A.1 Metrics
We quantitatively evaluate mesh quality by uniformly sampling 100K points from the faces of both the ground truth meshes and the predicted meshes, and then computing a set of metrics to assess various aspects of the reconstruction.
For mesh extraction, we report the following metrics: Chamfer Distance (CD) to evaluate the overall quality of a reconstructed mesh; Edge Chamfer Distance (ECD) to assess the preservation of sharp edges by sampling points near sharp edges and corners; and Normal Consistency (NC) to evaluate the quality of the surface normals. Additionally, we report the number of mesh vertices (#V) and the number of mesh faces (#F). We also provide the ratio of the estimated number of vertices to the ground truth number of vertices (#V_R) and the same ratio for faces (#F_R).
For mesh generation, Coverage (COV) captures the diversity of generated meshes and is sensitive to mode dropping, but it does not reflect the quality of the results. Minimum Matching Distance (MMD) measures the average distance between the reference set and their nearest neighbors in the generated set, though it lacks sensitivity to low-quality outputs. The 1-Nearest Neighbor Accuracy (1-NNA) assesses both quality and diversity between the generated and reference sets. To evaluate topology quality, we render the ground truth meshes and generated meshes with their wireframes visualized. We then employ Frechet Inception Distance (FID) and Kernel Inception Distance (KID) on rendered images. MMD, and KID scores are scaled by a factor of $10^{3}$.
A.2 Experiments
Additional Qualitative Experiments We present more qualitative results of MeshAnything here. As shown in Fig. 5 and Fig. 6, MeshAnything effectively generates AMs from various 3D representations. When integrated with different 3D assets production pipelines, our method effectively achieves mesh generation with diverse conditions.
Next, Fig. 6 demonstrates that MeshAnything does not simply overfit but understands how to generate meshes with efficient topology that conform to the given shape. To prove this, we use manually-created meshes as ground truth and use their shapes as conditions to test whether our model can generate meshes with comparable topology. To effectively use the ground truth as conditions, we first convert them into dense meshes using Marching Cubes Lorensen & Cline (1987) to disrupt their face structure. Then, we sample point clouds with normals from the dense meshes to serve as shape conditions. The experimental results in Fig. 6 show that MeshAnything is capable of generating meshes comparable to or even surpassing those modeled by human artists, exhibiting diverse and strong 3D modeling capabilities.
Comparison with mesh extraction baselines. Our method is related to various mesh extraction methods Lorensen & Cline (1987); Chen & Zhang (2021); Chen et al. (2022); Shen et al. (2023); Peng et al. (2021) since we also convert other 3D representations into meshes. However, it is important to note that previous approaches are reconstruction-like methods that produce dense meshes, while our approach is generative, creating Artist-Created Meshes (AMs) that are significantly more complex to produce than dense meshes. Therefore, strictly speaking, our method cannot be considered the same as these reconstruction-based mesh extraction methods. The main purpose of this comparison is to use these mesh extraction methods as a reference for evaluating the quality of the meshes generated by MeshAnything in terms of shape. We compare MeshAnything with Lorensen & Cline (1987); Shen et al. (2023); Peng et al. (2021). Among these, MarchingCubes is the most popular mesh extraction method, FlexiCubes represents the state-of-the-art in mesh extraction, and Shape as Points is the leading method for extracting mesh from point cloud.
We also combined these methods with the remesh technique to test whether they could significantly reduce the number of faces while maintaining shape quality. We used Blender Remesh in voxel mode Community (2018); Blender Development Team (2024), specifically using Blender version 4.1, as the remesh method. Since our evaluation dataset includes non-watertight meshes, we first extract the signed distance fields (SDF) of all ground truth meshes using Wang et al. (2022), which can handle non-watertight meshes. We then apply Marching Cubes with a resolution of 128 on these SDFs. Next, we apply Blender remesh Blender Development Team (2024) with different voxel sizes to the Marching Cubes results, as both the remesh method and our approach are capable of simplifying topology. Additionally, the Marching Cubes result is used as the shape condition input to MeshAnything to obtain our results. The settings of Shen et al. (2023) and Peng et al. (2021) follow their papers.
As shown in Tab. 5, we found that these methods require hundreds of times more faces to achieve results comparable to our method. Comparing (a), (g), (i) and (k), our method lags in Chamfer Distance (CD) and Normal Consistency (NC), mainly due to our method’s inherent failure cases as a generative model, which makes it less robust than these reconstruction-based mesh extraction methods. When comparing with remesh methods, we observe that they incur a high cost to achieve a face count similar to ours. Comparing (f) and (k), we find that even when remesh methods achieve a comparable face count, the number of vertices is still several times higher than ours, indicating that the topology efficiency of remesh methods is far inferior to ours, as they completely ignore the shape characteristics of the 3D assets. It’s important to note that the metrics in mesh etraction can only indicate the quality of shape alignment, which do not effectively reflect the topological advantages of our method. Additionally, we surprisingly find that our method can produce results with fewer faces than the ground truth, demonstrating that MeshAnything is not overfitting to the data but instead learns an efficient topology representation, occasionally surpassing the ground truth meshes.
@@ A.2 Experiments
Ablations on Noise-Resistant Conditional Decoder. We perform ablation experiments to verify the effectiveness of the Noise-Resistant Decoder. We begin with a VQ-VAE trained without any noise or conditioning. We then perform ablation between two settings: one where the decoder remains unchanged and unaware of the shape condition, and another where the shape condition is injected into the transformer, as described in Section 4.2. Next, we randomly sample a noise from gumbel distribution and add it to codebook sampling logits during the vector quantization process to simulate the potential low-quality token sequences generated by the transformer. We control the noise level by scaling the added noise.
Table 6: Experiments on the Impact of Input Point Cloud Quality on Generated Results.
[…]
(×$\times 10^{-2}$) (×$\times 10^{-2}$) (×$\times 10^{3}$) (×$\times 10^{3}$)
(a) Noise scale 0.005 2.351 6.412 0.897 0.175 0.321 0.895 0.880
(b) Noise scale 0.020 2.980 6.970 0.881 0.180 0.330 0.901 0.910
(c) Noise scale 0.050 4.910 8.556 0.755 0.162 0.284 0.811 0.802
(d) Rodin 2.552 6.622 0.833 0.185 0.342 0.919 0.923
(e) MeshAnything 2.256 6.245 0.902 0.172 0.318 0.888 0.871
@@
Experiments on the Impact of Input Point Cloud Quality on Generated Results. MeshAnything takes point clouds as input, and its robustness to point cloud quality determines its versatility across various applications. We design two experiments to evaluate its tolerance to input point cloud quality: First, keeping the other evaluation settings unchanged, we apply Gaussian noise to the input point cloud coordinates and normals. Specifically, for each point, we randomly sample Gaussian noise from a standard distribution, scale it by a noise factor, and add it to the point’s coordinates. The same approach is applied to the normals, but normalization is applied after adding the noise. Second, we use Rodin’s generation result as the ground truth mesh, sample point clouds from this mesh as input, and evaluate the deviation between the generated result and the ground truth.
As shown in Tab. 6, MeshAnything did not experience a significant performance drop in (a) and (b), demonstrating resilience to noise in the point cloud, with a noticeable performance decrease only in (c). It is important to note that the input point cloud is normalized to the range [-1,1], and the noise scale in (c) is already quite large. The experiment in (d) further demonstrates that MeshAnything can tolerate generated point clouds and effectively integrate with 3D generation models.
A.3 Limitations
Our method cannot generate meshes that exceed the maximum face count limit, so it cannot convert large scenes and particularly complex objects into meshes. Additionally, due to its generative nature, our method is not as stable as reconstruction-based mesh extraction methods like Lorensen & Cline (1987); Shen et al. (2023).
A.4 Social Impact